An epicyclic reduction gear, as shown in Fig. 13.38, has a shaft A fixed to arm B. The arm B has a pin fixed to its outer end and two gears C and E which are rigidly fixed, revolve on this pin. Gear C meshes with annular wheel D and gear E with pinion F. G is the driver pulley and D is kept stationary.


The number of teeth are : D = 80 ; C = 10 ; E = 24 and F = 18. If the pulley G runs at 200 r.p.m. ; find the speed of shaft A.


1
Expert's answer
2021-06-02T07:43:27-0400

Consider gear G , then "y+x=200 \implies y=200-x"

Gear D is staionary "So, y-\frac{-18+x}{24*8}=0 \implies 200-x=\frac{3}{32}x \implies x=182.857"

Shaft A is fixed to arm B

Arm B , N=y

NB =200-x

NB =200-182.857=17.14 rpm

Therefore, NB =NA =17.14 rpm